William Brown, the CFO of Oriole Automotive, Inc., is putting together this year's financial statements. He has gathered the following balance sheet information: The firm had a cash balance of $23,015, accounts payable of $163,257, common stock of $311,300, retained earnings of $512,159, inventory of $213,100, goodwill and other assets equal to $78,656, net plant and equipment of $710,100, and short-term notes payable of $21,115. It also had accounts receivable of $141,258 and other current assets of $11,223. How much long-term debt does Oriole Automotive have?

Long-term debt $________
Answer #1

Total asset = Cash + inventory+goodwill+PPE+receivables+current assets=23015+213100+78656+710100+141258+11223=1177352

LTD = total asset-accnt payable-common stock -retained earnings-short term notes = 1177352-163257-311300-512159-21115=

169521